OpenStudy (itiaax):

Find the least value of |z-1| if |z+3-i|=2

OpenStudy (gorv):

basically it is circle |z+3-i|=|z-(-3+i)|=2 with radius= 2 and centre at (-3,i)

OpenStudy (gorv):

|z-1| implies circle with centre at (1,0)

OpenStudy (gorv):

let z= x+iy |x+iy+3-i|=|x+3+i(y-1)|=2 now taking its magnitude \[\sqrt{(x+3)^{2}+(y-1)^{2}}=2\]

OpenStudy (gorv):

or \[(x+3){^2}+(y-1)^{2}=4\]

OpenStudy (gorv):

|z-1|=|x+iy-1| =\[\sqrt{(x-1){2}+y^2}\]

OpenStudy (gorv):

now convert first equation in the below given

OpenStudy (itiaax):

What equation am I converting?

OpenStudy (gorv):

(x+3)^2+(y-1)^2=4 to (x-1)^2+y^2

OpenStudy (itiaax):

I don't understand :S

OpenStudy (gorv):

we had the value of (x+3)^2+(y-1)^2 and need to find min of (x-1)^2+y^2

OpenStudy (gorv):